OpenStudy (anonymous):

Solve the problem. Where appropriate, round to the nearest tenth of a percent. The price of a printer was reduced from $400 to $180. What was the percent decrease? A. 60% B. 122.2% C. 45% D. 55%

OpenStudy (anonymous):

400-180=220 220/400= x/100 220*100=400x x=55 The answer is 55%